Consider two projects with the following cash flows: Project L (Cfo = -500, C01 = 200, C02 = 500, C03 = 700), Project S ( (Cfo = -300, C01 = 400, C02 = 200, C03 = 100). What is the cross-over rate for these projects?

45%

30%

20%

10%
Answer #1

Answer is 45%

Let crossover rate be i%

Project L:

NPV = -500 + 200/(1+i) + 500/(1+i)^2 + 700/(1+i)^3

Project S:

NPV = -300 + 400/(1+i) + 200/(1+i)^2 + 100/(1+i)^3

NPV of Project L = NPV of Project S at crossover rate
-500 + 200/(1+i) + 500/(1+i)^2 + 700/(1+i)^3 = -300 + 400/(1+i) + 200/(1+i)^2 + 100/(1+i)^3
-200 - 200/(1+i) + 300/(1+i)^2 + 600/(1+i)^3 = 0

Using financial calculator, i = 45%

Crossover Rate = 45%